OpenStudy (chycora):

3. (1 pt) What is the solution to the equation ? k =

OpenStudy (chycora):

i think it's 7k

OpenStudy (chycora):

@Nibby

OpenStudy (chycora):

@iGreen.

OpenStudy (anonymous):

No, I don't' think that's right.. What's -7 * 8?

OpenStudy (chycora):

-56

OpenStudy (chycora):

@iGreen.

OpenStudy (anonymous):

Yes, now add-56 + 10

OpenStudy (anonymous):

@chycora

OpenStudy (chycora):

k -46

OpenStudy (holly00d1248):

now divide -46/2 and youll get your K=

OpenStudy (holly00d1248):

@chycora

OpenStudy (holly00d1248):

make sense

OpenStudy (texaschic101):

(2k - 10) / -7 = 8 --- multiply both sides by -7 cancelling out the -7 on the right side 2k - 10 = -7 * 8 2k - 10 = -56 2k = -56 + 10 2k = -46 k = -46/2 k = - 23 check.. (2k - 10) / -7 = 8 (2(-23) - 10) / -7 = 8 (- 46 - 10) / -7 = 8 -56/-7 = 8 8 = 8 (correct) so k = -23 any questions ?

OpenStudy (texaschic101):

oops..cancelling out the -7 on the left side

OpenStudy (chycora):

-23